A production line operation is tested for filling weight accuracy using the following hypotheses.

Hypothesis Conclusion and Action
H 0:  = 16 Filling okay, keep running
H a:     16 Filling off standard; stop and adjust machine

The sample size is 39 and the population standard deviation is  = 0.9. Use  = .05. Do not round intermediate calculations.

  1. What would a Type II error mean in this situation?
    SelectConcluding that the mean filling weight is not 16 ounces when it actually isConcluding that the mean filling weight is 16 ounces when it actually isn'tItem 1  
  2. What is the probability of making a Type II error when the machine is overfilling by .5 ounces (to 4 decimals)?
      
  3. What is the power of the statistical test when the machine is overfilling by .5 ounces (to 4 decimals)?
      
  4. Choose the power curve for this hypothesis test
Answer #1

Given that :

n = 39

\sigma=0.9

\alpha= 0.05

Ho \mu = 16

Ha  \mu\neq 16

a) fail reject null hypothisis Ho, when Ho is false

population mean is asumed to equal to 16 actually different from 16

b) the critical value table corrosponding a probablity 0.025 / 0.975

z = \pm 1.96

z score of standred diviation

x = \mu + z(\sigma/n) = 16 - 1.96(0.8/\sqrt{39}) = 15.7836

x = \mu + z(\sigma/n) = 16 + 1.96(0.8/\sqrt{39}) = 16.2163

calculate z value by using standres diviation

z = x - \mu / (\sigma/n) = 15.7836 - 16.5 / (0.8 /\sqrt{39}) = -5.59

z = x - \mu / (\sigma/n) = 16.2163 - 16.5 / (0.8 /\sqrt{39}) = -2.21

probability using table

P = P( -5.59 < Z < -2.21) = P ( Z < -2.21 ) - ( Z < -5.59 ) = 0.0136 - 0 = 0.0136

c) the power of complement calculated part b

POWER = 1- 0.0136 0.9864

d)